# Bias Injection Attacks on RAG Databases and Sanitization Defenses

### v1

@@ 3 Bias Injection Attack on RAG
Workflow of bias injection attack: The core of bias injection is to produce human-like yet biased text of a target question. Because the attacker controls the content fully, we can use an LLM to help generate these passages. Prior work suggests that dense retrievers tend to favor LLM-written content in retreival, so we can expect that using a state-of-the-art LLM to produce text will be effective [35, 37]. The attack process for a target question follows several steps:
@@
  • (1) The attacker first needs $J$ passages with different perspectives on the question $q$, to serve as candidate passages for the attack. One approach is to retrieve these passages via keyword search using search engines. Alternatively, the attacker may prompt an LLM to generate them, using the Prompt-Synthetic in Appendix B as the prompt:
@@
  • (2) The attacker employs an embedding model $E_{d}(\cdot)$ to project the candidate passages into the embedding space and performs PCA [36] on the embeddings $E_{d}(d_{\text{synt},j}),j\in[1,J]$. By projecting the passage embeddings onto the resulting polarization axis, the attacker obtains the scalar PS value $\text{PS}(d_{\text{synt},j})$ for each passage, which captures its perspective bias.
@@
  • (3) The attacker selects a seed passage $d_{\text{seed}}$ by choosing the passage with the highest or lowest PS:
    $d_{\text{seed}}=\mathop{\arg\max}_{d_{\text{synt},j}}\text{PS}(d_{\text{synt},j})\text{ or }\mathop{\arg\min}_{d_{\text{synt},j}}\text{PS}(d_{\text{synt},j})$ (6)
@@
    aiming to induce a positive or negative shift in the generated answer along the polarization axis [^3] We note that the terms “positive” and “negative” are used solely to distinguish opposing perspectives (e.g., supporting vs. opposing a policy), and do not imply any evaluative or emotional judgment about the content of the viewpoints themselves.. For convenience, we refer to these as positive bias injection and negative bias injection, respectively.
@@
  • (4) The attacker then repeatedly queries the LLM to generate $n$ adversarial passages for the question $q$ that share the same perspective as the seed passage $d_{\text{seed}}$
    $d_{\text{adv},j}=\textbf{LLM}(\text{Prompt-Adv},q,d_{\text{seed}}),\quad j\in[1,n].$ (7)
@@
    To ensure that the generated passages are factually accurate and semantically coherent, we carefully design the prompt (see Prompt-Adv in Appendix B) to guide the model toward grounded and contextually relevant content. We also control decoding parameters—such as using a low temperature and a high nucleus sampling threshold [38]—to reduce randomness in generation, thereby limiting unwarranted speculation and improving alignment with $d_{\text{seed}}$.
@@
  • (5) Finally, the $k$ adversarial passages are injected into the database. Once retrieved, these passages introduce varying degrees of bias into the generation process.
Properties of adversarial passages generated by bias injection attacks: We evaluate the proposed attack on the WIKI-BALANCE [28] dataset (detailed in Sec. 4.2) when no protection is in place. For each query, we use JJ=8 LLM-generated synthetic passages as candidate inputs and execute the attack workflow to construct nn=10 adversarial passages. We then assess whether the generated adversarial passages satisfy the expected properties defined in Sec.2.3 by comparing them against relevant benign passages from the WIKI-BALANCE dataset.
Table I shows the distribution of adversarial passages generated through our workflow. Specifically, more than 74% of the adversarial passages, across both negative and positive bias injections, satisfy both Property 1 and Property 2 in Sec. 2.3. They exhibit higher SS values to the query compared to benign passages, as the LLM generates them by leveraging the semantic relevance between the query and the output content in the embedding space [35, 37]. In addition, they hold biased perspectives due to their similarity to the selected seed passage with extreme PS values. Finally, the generator is given instructions that enforce factual correctness and semantic coherence. This ensures that the adversarial passages are indistinguishable from benign content (Property 3) and can evade detection by automated fact-checking APIs (see more results in Sec. 4.5).
When these adversarial passages are retrieved as context for question answering, they cause the LLM to produce a biased output that diverges significantly from the answer generated using benign context (see the examples in Fig. 2). A more comprehensive evaluation within a real RAG system, where only retrieved adversarial passages are included in the context, is presented later in Sec. 4.
4 Prior Sanitization Defenses vs. Our Attack
@@ 4.1 Existing Sanitizers and Their Weaknesses
The most basic countermeasures against bias injection are methods that try to maximize relevance of retrieved passages to the query, i.e., the similarity score (SS). Along this direction, dense-retrieval methods, such as DPR [39] and Sentence-BERT [31], are widely adopted in RAG systems due to their effectiveness in capturing semantic similarity. However, these methods typically do not have any built-in defense mechanisms against knowledge corruption. Relying solely on query-passage relevance, they are particularly vulnerable to the proposed bias injection attack, as adversarial passages with high similarity can easily crowd out benign passages from the top-$k$ retrieval results—thereby influencing the perspective of the generated answers. We assume that our undefended baseline employs these mechanisms.
@@
Negative Bias Injection 74.7% 16.2% 7.3% 1.8%
Positive Bias Injection 78.0% 10.3% 8.7% 3.0%
TABLE I: Among the 4,520 adversarial passages generated by our workflow for 452 queries in WIKI-BALANCE [28], more than 74% satisfy both Property 1 and Property 2.
In contrast to these relevance-only retrievers, several perspective-aware retrieval methods have been proposed to balance relevance, diversity, and conflict among the retrieved passages. These methods exhibit higher robustness against database corruption and can serve as potential defense mechanisms for bias injection attacks. Prominent examples of these include MMR [25] and SMART [26]. MMR reduces redundancy by penalizing candidates that are highly similar to already selected passages. It iteratively selects passages based on a relevance-diversity score, which accounts for both similarity to the query and dissimilarity to previously selected passages, thereby encouraging content diversity. SMART employs a more sophisticated selection strategy to optimize the trade-off among relevance, inter-passage similarity, and conflict resolution. It first constructs a conflict matrix between passages using a Natural Language Inference (NLI) model, along with a passage similarity matrix and a query-context relevance matrix. These are combined to form a conflict-aware kernel matrix for sampling using Determinantal Point Processes [40]. SMART then repeatedly computes the determinants of submatrices to estimate the marginal utility of each candidate and select the final subset.
@@
Fig. 3(a) illustrates the iterative retrieval process shared by both MMR and SMART. The first passage selected is the one most similar to the query (i.e., highest SS), while the second is chosen to be both relevant and diverse relative to the first. This procedure continues sequentially until the top-$k$ passages are selected. Although these methods are effective in avoiding repetition and limiting the dominance of adversarial passages, they cannot fully eliminate adversarial content—particularly when such passages are themselves diverse in form.
Unlike the above methods, BRRA [24] introduces query perturbation to enhance diversity. It generates multiple noise-perturbed variants of a query and retrieves passages relevant to both the original and perturbed queries (illustrated as green circles in Fig. 3(b)). It then re-ranks all retrieved passages based on retrieval frequency and rank to promote coverage across the semantic space. BRRA is effective in low-dimensional embedding spaces, where noise-induced perturbations cause significant angular deviation. However, in high-dimensional settings, random noise tends to produce only minor directional shifts after normalization, due to the concentration of measure on the unit hypersphere [41]. As a result, the similarity between original and perturbed queries remains high (e.g., >>0.9), yielding minimal variation in the retrieved results.
4.2 Benchmark for Bias Injection Attack
We built an RAG benchmark to evaluate the effectiveness of existing methods against the bias injection attack.
Datasets: Our experiments are mainly based on an open-source dataset, WIKI-BALANCE [28], containing 452 open-ended questions and a corpus of 4662 real Wikipedia pages relevant to these questions. While using WIKI-BALANCE as the primary dataset, we also applied our attack and defense methods to Reddit-Dialogues [^4] Each query in this dataset includes a Reddit topic and previous responses of users, the RAG generates a new response based on retrieved passages from a given knowledge database. [30] and HotpotQA [^5] We use ChatGPT-5 to extend some close-form questions (e.g., ”Were Scott Derrickson and Ed Wood of the same nationality?”) to open-ended questions (”How did Scott Derrickson’s and Ed Wood’s cultural and national backgrounds shape their filmmaking styles and themes?”). [29] to show the generalization across datasets.
@@ 4.2 Benchmark for Bias Injection Attack
  • Retriever: We use msmarco-distilbert-base-tas-b [42] as the encoder for our dense retriever. This pre-trained model is fine-tuned on large-scale passage ranking datasets using contrastive learning and has demonstrated strong performance in dense retrieval tasks [31]. During retrieval, both queries and documents are encoded into dense vectors, and their relevance is computed via cosine similarity. The retriever first coarsely selects the top-$4\times k$ most relevant candidate passages from the knowledge database, and then applies existing methods to refine the selection to the final top-$k$ passages.
  • Generator: We deploy LLMs with various model sizes—including Meta-Llama-3-8B [43], DeepSeek-R1-Distill-Qwen-14B [44], and GPT-4.1 [2] (via the OpenAI API)—using the vLLM framework [45] to generate answers to open-ended questions. A prompt containing the query, the $k$ retrieved passages, and critical thinking instructions is constructed to initiate the answer generation process (see Prompt-Generation in Appendix B).
Attacker settings: As outlined in Sec. 3, the attacker deploys a DeepSeek-R1-Distill-Qwen-14B model to generate $J=8$ synthetic passages. These passages are then evaluated using a separate sentence-t5-xl [46] encoder to calculate their PS values, as the RAG’s own encoder is unknown to the attacker. The attacker selects two passages with the highest and lowest PS values as seed passages. Based on the seed passages, the attacker prompts the model to generate $n=10$ adversarial variants for the positive bias injection and another 10 for the negative bias injection and finally injects them into the target knowledge database.
@@
Except for the separated results on generalization across LLM models and datasets, we present only the results for the most representative setting: DeepSeek-R1-Distill-Qwen-14B [44] as the generator and WIKI-BALANCE [28], a typical dataset featuring open and controversial questions across diverse topics, as the evaluation dataset. The complete code, data, and results will be released in our artifact.
@@
  • Recall@k: The proportion of correctly identified relevant benign passages among the top-$k$ results, assessed using Qrels [^6] Qrels (manual relevance judgments) are standard evaluation resources that contain query-document pairs labeled with binary or graded relevance scores. These labels are annotated by assessors following strict guidelines and are widely used as ground truth for evaluating retrieval quality.. A higher retrieval effectiveness score implies that the retrieved context is more informative and aligned with the intended question.
  • Polarization Score (PS): As defined in Sec. 2.2, PS quantifies the semantic bias of a passage. PS shift—the magnitude of the difference between the PS values in the attacked and unattacked cases—measures the impact of attacks on the bias level of the content.
Compared baselines: We compare the potential baselines described in Sec. 4.1:
@@
4.3 Impact of Bias Injection Attacks on Retrieval
@@
TABLE II: Average A-Recall@5, the proportion of adversarial passages in top-5 retrieved passages, for varying number of attacker-injected passages.
We evaluate how bias injection attacks affect different retrievers by examining two key aspects on the top-5 passages retrieved by the retrievers: A-Recall@5 and Recall@5. An effective defense should minimize A-Recall@5 as much as possible, without sacrificing the overall retrieval quality measured by Recall@5.
A-Recall@5: Table II presents the average A-Recall@5 across all 452 WIKI-BALANCE queries under different injection intensities, where “#Positive Injection” and “#Negative Injection” denote the proportions of injected positive and negative adversarial passages respectively. As the injection intensity increases, all retrievers exhibit higher A-Recall@5. Among the baselines, the No Def. baseline demonstrates the highest vulnerability to bias injection attacks, attributable to its exclusive reliance on query-passage similarity for ranking. In contrast, MMR, SMART, and BRRA incorporate diversity criteria (Sec. 4.1) that partially mitigate adversarial retrieval, yet still leave over $19\%$ of adversarial passages unsanitized at injection intensities 5 and 10. This is nearly a $5\times$ reduction compared to No Def., but it comes at the cost of reduced Recall@5 (see Table III).
@@
TABLE III: Average Recall@5, the proportion of relevant passages (as labeled by Qrels) among the top-5 retrieved passages. The $\pm$ values for ”Unattacked” indicate the 99% confidence interval, quantifying the sampling noise.
@@
Worst-case performance: Since the attacker can vary the number of injected adversarial passages, we compute the maximum ratio between A-Recall@5 between Recall@5 across all injection intensities to assess the worst-case performance of each defense method. Specifically, the maximum ratios are as follows: 98 (at #Neg=10) for No Def., 0.85 (at #Neg=1,5 and #Pos=5,10) for MMR, 0.86 (at #Neg=1) for SMART, and 20 (at #Neg=10) for BRRA. These results demonstrate that both No Def. and BRRA are highly susceptible to adversarial passage domination, with adversarial content frequently overwhelming the top-5 retrievals. In contrast, MMR and SMART reduce this vulnerability; however, they still allow a number of unsanitized adversarial passages that is comparable to the number of relevant benign passages.
@@
TABLE IV: Average PS shift of the top-5 retrieved passages, expressed as a percentage of the unattacked Avg. $|\text{PS}|$.
@@
TABLE V: Average PS shift of answers generated by DeepSeek-R1-Distill-Qwen-14B [44], expressed as a percentage of unattacked $|PS|$. The answers show a significant PS shift due to the biased context passages retrieved.

### v2

@@ 4 EBI Attack on RAG Databases
Workflow of EBI attack: The core of EBI is to produce human-like yet epistemically biased text for a target question. Because the attacker controls the content fully, we can use an LLM to help generate these passages. Prior work suggests that dense retrievers tend to favor LLM-written content in retrieval [31, 33], so we can expect that using a state-of-the-art LLM to produce text will be effective. The attack process for a target question follows several steps:
@@
  • (1) The attacker first needs $J$ passages with different perspectives on the question $q$, to serve as candidate passages for the attack. One approach is to retrieve these passages via keyword search using search engines. In practice, the attacker may prompt an LLM to generate them, using the Prompt-Synthetic in Appendix A as the prompt:
@@
  • (2) The attacker employs an encoder $E(\cdot)$ to project the candidate passages into the embedding space and performs PCA [30] on the embeddings $\{E(d_{\text{synt},j})\}_{j\in[1,J]}$, without requiring any viewpoint labels. In other words, these candidate passages also serve as the anchor passages to derive the polarization axis. By projecting the passage embeddings onto the resulting polarization axis, the attacker obtains their PS values $\{\text{PS}(d_{\text{synt},j})\}_{j\in[1,J]}$.
@@
  • (3) The attacker selects the most biased passage $d_{\text{bias}}$ by choosing the passage with the highest or lowest PS:
    $d_{\text{bias}}=\mathop{\arg\max}_{d_{\text{synt},j}}\text{PS}(d_{\text{synt},j})\text{ or }\mathop{\arg\min}_{d_{\text{synt},j}}\text{PS}(d_{\text{synt},j})$ (7)
@@
    aiming to induce a positive or negative shift in the generated answer along the polarization axis [^3] We note that the terms “positive” and “negative” are used solely to distinguish opposing viewpoints (e.g., supporting vs. opposing a policy), and do not imply any evaluative or emotional judgment about the content of the viewpoints themselves.. For convenience, we refer to these as positive EBI and negative EBI, respectively.
@@
  • (4) The attacker then repeatedly queries the LLM to generate $n$ adversarial passages for the question $q$ that mimic the viewpoint of $d_{\text{bias}}$ while varying their presentation
    $d_{\text{adv},j}=\textbf{LLM}(\text{Prompt-Adv},q,d_{\text{bias}}),\quad j\in[1,n].$ (8)
@@
    To ensure that the generated passages are factually accurate and semantically coherent, we carefully design the prompt (see Prompt-Adv in Appendix A) to guide the model toward grounded and contextually relevant content. We also control decoding parameters—such as using a low temperature and a high nucleus sampling threshold [34]—to reduce randomness in generation, thereby limiting unwarranted speculation and improving alignment with $d_{\text{bias}}$.
@@
  • (5) Finally, the $n$ adversarial passages are injected into the database. Once retrieved, these passages introduce attack-desired epistemic bias into the generation process.
Satisfying attack conditions in practice: We explicitly design the attack workflow to satisfy each of the three attack conditions defined in Sec. 3. To satisfy,
  • Property 1 (High relevance), adversarial passages are generated in a query-conditioned manner, prompting the LLM to produce content that directly addresses the target question. This design encourages strong semantic alignment between the generated passages and the query in the embedding space [31, 33].
  • Property 2 (Epistemic bias), we guide generation using a biased passage selected from an extreme end of the polarization axis, thereby steering the LLM to consistently express a targeted viewpoint. Table 1 shows a property check result from our benchmark detailed in Sec. 5.2, which validates the effectiveness of our attack.
  • Property 3 (Stealth), we instruct the LLM to maintain factual correctness and linguistic coherence, avoiding explicit signals such as false statements or unnatural phrasing that could serve as detectable fingerprints.
5 Prior Sanitization Defenses vs. EBI Attack
@@ 5.1 Existing Sanitizers and Their Weaknesses
The most basic countermeasures against EBI to consider are those which rely solely on maximizing the relevance (i.e., SS value) of the retrieved passages. Along this line of defenses, dense-retrieval methods such as DPR [35] and Sentence-BERT [26], are widely adopted in RAG systems due to their effectiveness in capturing semantic similarity. However, they typically do not have any built-in defense mechanisms against EBI. Because retrieval decisions are driven solely by relevance, adversarial passages that are both biased and highly similar to the query can be easily retrieved. Our undefended baseline employs these mechanisms by default.
@@
Negative EBI 74.7% 16.2% 7.3% 1.8%
Positive EBI 78.0% 10.3% 8.7% 3.0%
Table 1: Among the 4,520 adversarial passages generated by our workflow for 452 queries in WIKI-BALANCE [20], more than 74% satisfy both Property 1 and Property 2.
In contrast to these relevance-only retrievers, several perspective-aware retrieval methods have been proposed to balance relevance, diversity, and conflict among the retrieved passages. They can serve as potential defense mechanisms for EBI attacks. Prominent examples of these include MMR [24] and SMART [25]. MMR reduces redundancy by penalizing candidates that are highly similar to already selected passages. It iteratively selects passages based on a relevance-diversity score, which accounts for both similarity to the query and dissimilarity to previously selected passages, thereby encouraging content diversity. SMART employs a more sophisticated selection strategy to optimize the trade-off among relevance, inter-passage similarity, and conflict resolution. Although these methods are effective in avoiding repetition and limiting the dominance of adversarial passages, as we will show later, they cannot fully eliminate EBI adversarial content—the diversity mechanism inadvertently encourages the retrieval of both benign and adversarial passages (see Fig. 3(a)).
@@
Lastly, unlike the above methods, BRRA [23] introduces query perturbation to enhance diversity. It generates multiple noisy perturbations to the query and retrieves passages relevant to both the original and perturbed queries (illustrated as green circles in Fig. 3(b)). It then re-ranks all retrieved passages based on retrieval frequency and rank to promote coverage across the semantic space. BRRA is effective in low-dimensional embedding spaces, where noise-induced perturbations cause significant angular deviation. However, in high-dimensional settings, random noise induces only minor directional shifts after normalization, due to the concentration of measure on the unit hypersphere [36]. As a result, the similarity between original and perturbed queries remains high (e.g., >>0.9), yielding minimal variation in retrieved results.
5.2 Benchmark for Measuring EBI Attacks
We built an RAG benchmark to evaluate the effectiveness of existing methods against our proposed EBI attack.
Datasets: Our experiments are mainly based on an open-source dataset, WIKI-BALANCE [20], containing 452 open-ended questions and a corpus of 4662 real Wikipedia pages relevant to these questions. While using WIKI-BALANCE as the primary dataset, we also applied our attack and defense methods to Reddit-Dialogues [^4] Each query in this dataset includes a Reddit topic and previous responses of users, the RAG generates a new response based on retrieved passages from a given knowledge database. [22] and HotpotQA [^5] We use ChatGPT-5 to extend some close-form questions (e.g., ”Were Scott Derrickson and Ed Wood of the same nationality?”) to open-ended questions (”How did Scott Derrickson’s and Ed Wood’s cultural and national backgrounds shape their filmmaking styles and themes?”). [21] to show the generalization across multiple datasets.
@@ 5.2 Benchmark for Measuring EBI Attacks
  • Retriever: We use msmarco-distilbert-base-tas-b [37] as the encoder for our dense retriever. This pre-trained model is fine-tuned on large-scale passage ranking datasets using contrastive learning and has demonstrated strong performance in dense retrieval tasks [26]. During retrieval, both queries and documents are encoded into dense vectors, and their relevance is computed via cosine similarity. The retriever first coarsely selects the top-$4\times k$ most relevant candidate passages from the knowledge database, and then applies existing methods to filter the final top-$k$ passages.
  • Generator: We deploy LLMs with various model sizes—including Meta-Llama-3-8B [38], DeepSeek-R1-Distill-Qwen-14B [39], and GPT-4.1 [2] (via the OpenAI API)—using the vLLM framework [40] to generate answers to open-ended questions. A prompt containing the query, the $k$ retrieved passages, and critical thinking instructions is constructed to initiate the answer generation process (see Prompt-Generation in Appendix A).
Attacker settings: As outlined in Sec. 4, the attacker deploys a DeepSeek-R1-Distill-Qwen-14B model to generate $J=8$ synthetic passages. These passages are then evaluated using a separate sentence-t5-xl [41] encoder to calculate their PS values, as the RAG’s own encoder is unknown to the attacker. The attacker selects two passages with the highest and lowest PS values as biased passages. Based on the biased passages, the attacker prompts the model to generate $n=10$ adversarial variants for positive EBI and another 10 for negative EBI and finally injects them into the knowledge database.
@@
Apart from the results about generalization across LLM models and datasets, we present attack efficacy results only for the most representative setting: DeepSeek-R1-Distill-Qwen-14B [39] as the generator and WIKI-BALANCE as the evaluation dataset [20]. This is a typical dataset featuring open and controversial questions across diverse topics. The complete code and data are released in our artifact.
@@
  • Recall@k: The proportion of correctly identified relevant benign passages among the top-$k$ results, assessed using Qrels [^6] Qrels (manual relevance judgments) are standard evaluation resources that contain query-document pairs labeled with binary or graded relevance scores. These labels are annotated by assessors following strict guidelines and are widely used as ground truth for evaluating retrieval quality.. A higher Recall@k implies that the retrieved context is more informative and aligned with the intended question.
  • Polarization Score (PS): As defined in Sec. 2.2, PS quantifies the viewpoint of a passage. PS shift—the magnitude of the difference between the PS values in the attacked and unattacked cases—measures the impact of attacks on the bias level of the content.
Compared baselines: We compare the potential baselines, including the defenses described in Sec. 5.1:
@@
5.3 Retrieval Performance under EBI Attacks
@@
Table 2: Average A-Recall@5, the proportion of adversarial passages in top-5 retrieved passages, for varying number of attacker-injected passages.
We evaluate how EBI attacks affect different retrievers by examining two key aspects on the top-5 passages retrieved by the retrievers: A-Recall@5 and Recall@5. An effective defense should minimize A-Recall@5 as much as possible, without sacrificing the retrieval quality measured by Recall@5.
A-Recall@5: Table 2 presents the average A-Recall@5 across all 452 WIKI-BALANCE queries under different injection intensities, where “#Positive Injection” and “#Negative Injection” denote the numbers of injected adversarial passages for positive EBI and negative EBI respectively. As the injection intensity increases, all retrievers exhibit higher A-Recall@5. Among the baselines, the No Def. baseline demonstrates the highest vulnerability to EBI attacks, attributable to its exclusive reliance on query-passage similarity for ranking. In contrast, MMR, SMART, and BRRA incorporate diversity criteria (Sec. 5.1) that partially mitigate adversarial retrieval, yet still leave over $19\%$ of adversarial passages unsanitized at injection intensities 5 and 10. This is nearly a $5\times$ reduction compared to No Def., but it comes at the cost of reduced Recall@5 (see Table 3).
@@
Table 3: Average Recall@5, the proportion of relevant passages (as labeled by Qrels) among the top-5 retrieved passages. The $\pm$ values for "Unattacked" indicate the 99% confidence interval, quantifying the sampling noise.
@@
Worst-case performance: Since the attacker can vary the number of injected adversarial passages, we compute the maximum ratio between A-Recall@5 and Recall@5 across all injection intensities to assess the worst-case performance of each defense method. Specifically, the maximum ratios are as follows: 98 (at #Neg=10) for No Def., 0.85 (at #Neg=1,5 and #Pos=5,10) for MMR, 0.86 (at #Neg=1) for SMART, and 20 (at #Neg=10) for BRRA. These results demonstrate that both No Def. and BRRA are highly susceptible to adversarial passage domination, with adversarial content frequently overwhelming the top-5 retrievals. In contrast, MMR and SMART reduce this vulnerability; however, they still allow a number of unsanitized adversarial passages that is comparable to the number of relevant benign passages.
@@
Table 4: Average PS shift of the top-5 retrieved passages, expressed as a percentage of the unattacked Avg. $|\text{PS}|$.
@@
Table 5: Average PS shift of answers generated by DeepSeek-R1-Distill-Qwen-14B [39], expressed as a percentage of unattacked average $|PS|$. The answers show a significant PS shift due to the biased context passages retrieved.